OpenStudy (shealyn):

What is the equation of the line which includes points (2, 6) and (0, 0)?

OpenStudy (joshtray):

Equation of a line is\[\frac{ y-y _{1} }{ x-x _{1} }=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (joshtray):

\[y _{1}=6 \] \[y _{2}=0\] \[x _{1}=2\] \[x _{2}=0\] Plug them in and solve